OpenStudy (anonymous):

Find the area bounded by the parabola y=x^2 and the line y=x-2

OpenStudy (saifoo.khan):

Best way is to try graphing if you can..

OpenStudy (anonymous):

THat's the problem... I don't know how to graph

OpenStudy (anonymous):

ok first you need to find point of intersections just set both equations equal x^2=x-2 x^2-x+2=0 @migo24 check your question because i am getting complex roots

OpenStudy (anonymous):

is the second equation y=x-2 or y=x+2 ??

OpenStudy (saifoo.khan):

Wait. lines aren't meeting. o_O

OpenStudy (anonymous):

yes thats what i am needing clarification .

OpenStudy (saifoo.khan):